Solve the system of equations by any method.
7x + 15y = 32
x = 3y + 20

Answer by mmr(:(1)

7x + 15y = 32
x = 3y + 20
you put x in for x.
7(3y+20)+15y=32 multiple by using the distrutive property.
21y+140+15y=32 add a like terms.
36y+140=32 take the opposite of 140(-140) and subtract it from both sides.
36y=-108 divide 36 to both sides.
y=-3
then you fill the y in
x=3y+20
x=3(-3)+20 multiple 3 times -3
x=-9+20 add -9 to 20
x=11
(11,-3)